When a pair of dice are rolled there are 36 different possible outcomes. If a pair of dice are rolled 3 times, what is the probability of getting a sum of 3 every time?

Question options:

a)

0.00017147

b)

0.0005787

c)

0.1111

d)

0.03703704
Answer #1

Answer

A pair of dice has 36 different possible outcomes

We have to find the probability of getting a sum of 3

So, possible events are (1,2) and (2,1) , i.e. only 2 possible outcomes

So, required probability = (favorable outcome)/(total outcome) = 2/36 = 1/18

Now, we have to find it for 3 rolls, so we can write it as

Probability =

where p = 1/18 and n= 3

setting the values, we get

Probability =

So, option A is correct answer